OpenStudy (anonymous):

The velocity given by v=2t^2-12t+8. find acceleration of particle after 4 sec, and find the minimum velocity?

OpenStudy (tkhunny):

Have you considered dv/dt = a(t), the acceleration?

OpenStudy (isaiah.feynman):

Find the derivative of the velocity function, then evaluate it at 4.

OpenStudy (anonymous):

I founded but the question ask about minimum Velocity? how to find it?

OpenStudy (anonymous):

how did u get it

OpenStudy (anonymous):

you plug in zero in which equation??? and how to find max velocity?

OpenStudy (loser66):

I don't think so

OpenStudy (loser66):

the function of the velocity is v = 2t^2 -12t+8 which is an upward parabola. Therefore, the minimum velocity is at its vertex, which is t = -b /2a = -(-12)/2*2 = 3.

OpenStudy (anonymous):

ok, this is t, but then plug in 3 which equation??

OpenStudy (loser66):

yes.

OpenStudy (loser66):

plug back to the original one

OpenStudy (loser66):

v = 2 t^2..........

OpenStudy (anonymous):

will get -10,, is that right!

OpenStudy (anonymous):

how about the max velocity!

OpenStudy (loser66):

I don't know how to get max if I don't have interval of time.

OpenStudy (anonymous):

what I understood, I should plug in zero in velocity equation!

OpenStudy (anonymous):

to get max velocity!!

OpenStudy (loser66):

https://www.desmos.com/calculator this is your graph, show me how to get the max?

OpenStudy (anonymous):

I don't know how to graph?

OpenStudy (loser66):

just do as I did, go to that site, type your function from the left hand side box and it automatically draw it for you. Of course, It is just for confirm our calculation.

OpenStudy (tkhunny):

There is no Maximum Velocity. The velocity function is unbounded. Please stop looking for the maximum velocity. You should know this from the nature of the parabola. That's why you studied Analytic Geometry prior to studying the calculus. Please listen to Loser66. Without a specific and finite interval of time, there is no maximum velocity.

OpenStudy (isaiah.feynman):

If you see the velocity function, it is the equation of a parabola opening up.